OpenStudy (henrietepurina):

Iris has been studying an invasive population of snails. This particular snail has no local predators so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years

OpenStudy (henrietepurina):

Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences.

OpenStudy (anonymous):

Which do you need help on?

OpenStudy (henrietepurina):

My reply!

OpenStudy (anonymous):

Ok well I just did this and for this pare you just form a function in terms of f(x) to model the population it can be as you choose.

OpenStudy (henrietepurina):

how do you do it?

OpenStudy (henrietepurina):

is still do understand the y intercept @pierodog

OpenStudy (henrietepurina):

*I* not is

OpenStudy (anonymous):

This question doesn't ask for y-intercept yet.

OpenStudy (henrietepurina):

Oh so can you pls help me from the equation

OpenStudy (henrietepurina):

*form* not from

OpenStudy (anonymous):

Well it really is up to you what do you want the number of snails in the beginning to be?

OpenStudy (henrietepurina):

i dunno 0?

OpenStudy (anonymous):

No a real number anything higher than 0.

OpenStudy (henrietepurina):

then 1

OpenStudy (henrietepurina):

can we finish please? @pierodog

OpenStudy (anonymous):

Ok so now use this function \[f(x)=a(b)^x\]

OpenStudy (anonymous):

a will be the principle meaning how much you start with.

OpenStudy (henrietepurina):

f(x) = 1(b)x

OpenStudy (anonymous):

Alrght, now b will be by how much it gets bigger, like if it doubles it would be 2, if it triples it would be 3, and so on and so forth.

OpenStudy (henrietepurina):

i dont know it! (B)

OpenStudy (anonymous):

You create this formula, it increases yearly by whatever you want, it could double, triple, quadruple, quintuple, whatever.

OpenStudy (henrietepurina):

can i do double?

OpenStudy (henrietepurina):

so umm... f(x) = 1(2)x

OpenStudy (anonymous):

Yes, I did.

OpenStudy (anonymous):

Alright, now x will equal the time, meaning how long it will be, I think the question asks for 15?

OpenStudy (henrietepurina):

so umm... f(x) = 1(2)15

OpenStudy (henrietepurina):

and the answer is 32768 if they start with 1!

OpenStudy (anonymous):

Yes, that would be your function

OpenStudy (henrietepurina):

OMG thanks a lot bud!

OpenStudy (anonymous):

You're welcome, I'm glad I could help.

OpenStudy (henrietepurina):

Now I get also how you do it!

OpenStudy (anonymous):

See? It's quite simple.